# Sample then Identify: A General Framework for Risk Control and Assessment in Multimodal Large Language Models

### v1

@@ 3.1 Preliminaries
Following standard SPC for risk control (Angelopoulos and Bates, 2021; Angelopoulos et al., 2024), we assume a calibration dataset ${\mathcal{D}}_{cal}=\left\{\left(X_{i},Y_{i}\right)\right\}_{i=1}^{N}$ of $N$ query-answer pairs and an independent and identically distributed (i.i.d.) [^2] Split conformal prediction requires that the test data points are drawn from the same distribution as the calibration data points. test point $\left(x_{test},y_{test}\right)$. Our goal of risk control is to construct prediction sets that are calibrated to bound the correctness miscoverage rate
@@ 3.2 Conformal Risk Control
A complete proof of Eq. 3 is presented in Appendix A.1.
@@
Inspired by self-consistency theory (Wang et al., 2022; Li et al., 2022a; Wang et al., 2024c), which states that a repetitively sampled response is viewed as a form of consistency linked to higher confidence in the response, we perform a simple semantic clustering process on the candidate set and obtain the frequency of each response, denoted as $F\left(\hat{y}_{m}^{(i)}\right)$. For a detailed description of semantic clustering, refer to Appendix B.1. Then, we define the nonconformity score of the $i$-th calibration data point as one minus the frequency of the acceptable response, denoted as $\hat{y}_{ref}^{(i)}$, within the candidate set, which is semantically equivalent to the label $Y_{i}$ (i.e., $\hat{y}_{ref}^{(i)}\Leftrightarrow Y_{i}$). $A\Leftrightarrow B$ represents a bidirectional entailment (MacCartney and Manning, 2015; Mohri and Hashimoto, 2024; Farquhar et al., 2024) between A and B (i.e., equivalence). The specific evaluation for bidirectional entailment is provided in Appendix B.1. In this case, the nonconformity score is formulated as
@@
where $\varepsilon=\alpha+\beta-\alpha\beta$. The complete proof of Eq. 6 is given in Appendix A.2.
@@ 4.1 Experimental Settings
We consider two multiple-choice VideoQA datasets: Video-MME (VMME) (Fu et al., 2024) and NExT-QA (NEXT) (Xiao et al., 2021), where we only utilize the multiple-choice section of NEXT, and two open-ended VideoQA datasets: MUSIC-AVQA (MUSC) (Li et al., 2022b) and MSVD (Chen and Dolan, 2011). More details of the dataset utilization can be found in Appendix B.
@@
We utilize the Empirical Error Rate (EER) to assess whether we control the risk of miscoverage by the two user-specified risk levels (Angelopoulos and Bates, 2021). Note that in the first stage, miscoverage refers to not sampling the acceptable responses, and this EER is controlled by $\alpha$. In the second stage, miscoverage refers to the calibrated prediction sets failing to cover the acceptable responses, and this EER is bounded by both $\alpha$ and $\beta$. Additionally, we adopt the commonly used Accuracy (ACC) and the Average Prediction Set Size (APSS) (Kostumov et al., 2024; Ye et al., 2024) to evaluate MLLM predictions. More details can be found in Appendix E.
@@ 4.2 Results for Risk Control
Following prior work (Mohri and Hashimoto, 2024; Su et al., 2024; Quach et al., 2024), we randomly split our datasets in half by default, calibrating the sampling number and threshold for identifying high-quality responses on the first half, i.e., the calibration set, and measure the EER (i.e., miscoverage rate) at various risk levels (i.e., upper bound) in two steps on the second half, i.e., the test set. We plot the empirical results in the two steps utilizing the GPT-4o mini model on four datasets in Figure 2(a) and Figure 2(b), and defer similar plots employing other MLLMs to Appendix F.
@@ 4.3 Results for Risk Assessment
In this section, we investigate the impact of semantic redundancy in the prediction set when employing the APSS metric to evaluate MLLMs, to address the research gap in open-ended settings. We fix the risk level in the first stage at 0.1 (i.e., $\alpha=0.1$) and measure the APSSs for different values of $\varepsilon$ ($=\alpha+\beta-\alpha\beta$), utilizing four open-source MLLMs on the VMME dataset. As demonstrated in Figure 4, the original APSS is sensitive to the error rate, leading to ambiguity in the evaluation of MLLMs at different risk levels, while after removing semantically redundant responses, the final APSS can consistently decrease as the error rate lowers, and accurately distinguish between foundational models of varying performance at different risk levels. Additionally, the post-processed prediction sets result in more efficient decision-making. Note that we still utilize bidirectional entailment detailed in Appendix B to determine the semantic equivalence between two responses.
@@ 5 Conclusion
Appendix A Proofs
A.1 Proof of equation 3
@@ A.2 Proof of equation 6
Appendix B Implementation Specifics
B.1 Semantic Clustering Process
@@ Appendix D Hyperparameters
Appendix E Evaluation Metrics
@@
Appendix F Additional Experimental Analysis

### v2

@@ 3.1 Preliminaries
Following standard SCP for risk control (Angelopoulos and Bates, 2021; Angelopoulos et al., 2024), we assume a calibration dataset ${\mathcal{D}}_{cal}=\left\{\left(X_{i},Y_{i}\right)\right\}_{i=1}^{N}$ of $N$ query-answer pairs and an independent and identically distributed (i.i.d.) [^2] Split conformal prediction requires that the test data points are drawn from the same distribution as the calibration data points. test point $\left(x_{test},y_{test}\right)$. Our goal of risk control is to construct prediction sets that are calibrated to bound the correctness miscoverage rate
@@ 3.2 Conformal Risk Control
A complete proof of Eq. 3 is presented in Appendix B.1.
@@
Inspired by self-consistency theory (Wang et al., 2022; Li et al., 2022a; Wang et al., 2024c), which states that a repetitively sampled response is viewed as a form of consistency linked to higher confidence in the response, we perform a simple semantic clustering process on the candidate set and obtain the frequency of each response, denoted as $F\left(\hat{y}_{m}^{(i)}\right)$. For a detailed description of semantic clustering, refer to Appendix C.1. Then, we define the nonconformity score of the $i$-th calibration data point as one minus the frequency of the acceptable response, denoted as $\hat{y}_{ref}^{(i)}$, within the candidate set, which is semantically equivalent to the label $Y_{i}$ (i.e., $\hat{y}_{ref}^{(i)}\Leftrightarrow Y_{i}$). $A\Leftrightarrow B$ represents a bidirectional entailment (MacCartney and Manning, 2015; Mohri and Hashimoto, 2024; Farquhar et al., 2024) between A and B (i.e., equivalence). The specific evaluation for bidirectional entailment is provided in Appendix C.1. In this case, the nonconformity score is formulated as
@@
where $\varepsilon=\alpha+\beta-\alpha\beta$. The complete proof of Eq. 6 is given in Appendix B.2.
@@ 4.1 Experimental Settings
We consider two multiple-choice VideoQA datasets: Video-MME (VMME) (Fu et al., 2024) and NExT-QA (NEXT) (Xiao et al., 2021), where we only utilize the multiple-choice section of NEXT, and two open-ended VideoQA datasets: MUSIC-AVQA (MUSC) (Li et al., 2022b) and MSVD (Chen and Dolan, 2011). More details of the dataset utilization can be found in Appendix C.
@@
We utilize the Empirical Error Rate (EER) to assess whether we control the risk of miscoverage by the two user-specified risk levels (Angelopoulos and Bates, 2021). Note that in the first stage, miscoverage refers to not sampling the acceptable responses, and this EER is controlled by $\alpha$. In the second stage, miscoverage refers to the calibrated prediction sets failing to cover the acceptable responses, and this EER is bounded by both $\alpha$ and $\beta$. Additionally, we adopt the commonly used Accuracy (ACC) and the Average Prediction Set Size (APSS) (Kostumov et al., 2024; Ye et al., 2024) to evaluate MLLM predictions. More details can be found in Appendix F.
@@ 4.2 Results for Risk Control
Following prior work (Mohri and Hashimoto, 2024; Su et al., 2024; Quach et al., 2024), we randomly split our datasets in half by default, calibrating the sampling number and threshold for identifying high-quality responses on the first half, i.e., the calibration set, and measure the EER (i.e., miscoverage rate) at various risk levels (i.e., upper bound) in two steps on the second half, i.e., the test set. We plot the empirical results in the two steps utilizing the GPT-4o mini model on four datasets in Figure 2(a) and Figure 2(b), and defer similar plots employing other MLLMs to Appendix G.
@@ 4.3 Results for Risk Assessment
In this section, we investigate the impact of semantic redundancy in the prediction set when employing the APSS metric to evaluate MLLMs, to address the research gap in open-ended settings. We fix the risk level in the first stage at 0.1 (i.e., $\alpha=0.1$) and measure the APSSs for different values of $\varepsilon$ ($=\alpha+\beta-\alpha\beta$), utilizing four open-source MLLMs on the VMME dataset. As demonstrated in Figure 4, the original APSS is sensitive to the error rate, leading to ambiguity in the evaluation of MLLMs at different risk levels, while after removing semantically redundant responses, the final APSS can consistently decrease as the error rate lowers, and accurately distinguish between foundational models of varying performance at different risk levels. Additionally, the post-processed prediction sets result in more efficient decision-making. Note that we still utilize bidirectional entailment detailed in Appendix C to determine the semantic equivalence between two responses.
@@ 4.4 Sensitivity Analyses
Figure 6: Stratified miscoverage rate at each size of prediction set on four VideoQA datasets utilizing the Gemini-1.5-Pro model. The solid line corresponds to Frequency, while the dashed line corresponds to Semantic Diversity.
Conditional Miscoverage Rate.
As mentioned before, TRON bounds the correctness miscoverage rate marginally over the test set (i.e., on average). However, for some critical applications requiring more stringent guarantees, users demand that the error rate is bounded for a particular test data point. In other words, the risk level should be guaranteed for any subset of the data points. In the most general case, conditional coverage is impossible to achieve (Angelopoulos and Bates, 2021). Following prior work (Kumar et al., 2023; Su et al., 2024), we set $\alpha=\beta=0.1$ (i.e., $\varepsilon=0.19$) and evaluate the size-stratified miscoverage rate utilizing the Gemini-1.5-Pro model. As demonstrated in Figure 6, the miscoverage rate fluctuates up and down with the set size, with some values exceeding $\varepsilon$. It is worth noting that when we employ Semantic Diversity as the reliability measurement, TRON’s conditional performance improves (e.g., the max miscoverage rate decreases from 0.2065 to 0.1667 on the VMME dataset). At this point, it’s like a bag of tricks, and we can enhance the capability of reliability measurement to optimize the conditional performance of risk control.
@@ 5 Conclusion
Appendix A Background of Conformal Prediction
Conformal prediction (CP) is model-agnostic and distribution-free, and can offer statistically rigorous guarantees of ground-truth coverage on fresh test samples with minimal assumptions (e.g., data exchangeability) given a relatively small calibration set [Angelopoulos and Bates, 2021, Campos et al., 2024]. CP can transform any heuristic notion of uncertainty from any model into a statistically rigorous one by constructing a prediction set adaptively to each particular input.
Let’s illustrate the framework utilizing classification problems. In the calibration set, $\left\{\left(X_{i},Y_{i}^{*}\right)\right\}_{i=1}^{N}$, we define the nonconformity score (NS), $s_{i}$, to identify a heuristic notion of the uncertainty of the decision process of the model for the $i$-th calibration data. Here, we utilize one minus the softmax output corresponding to the true class to link the NS with the uncertainty condition of correctness closely. Then we calculate the $\frac{\left\lceil\left(n+1\right)\left(1-\alpha\right)\right\rceil}{n}$ (i.e., approximate $1-\alpha$) quantile of all NSs (i.e., $\left\{s_{1},...,s_{n}\right\}$) in the calibration set and denote it as $\hat{q}$. Since there are approximately $1-\alpha$ calibration samples whose uncertainty states corresponding to their true class satisfy (i.e., $\leq$) the threshold $\hat{q}$, we construct the prediction set of each test sample following $\mathcal{C}\left({X}_{test}\right)=\left\{y\in[K]:s\left({X}_{test},y\right)\leq\hat{q}\right\}$. We expect that if the class $y$ satisfies $s\left({X}_{test},y\right)\leq\hat{q}$ (i.e., the uncertainty state of the current class $y$ satisfies the uncertainty condition corresponding to the true labels of approximate $1-\alpha$ calibration samples), $y$ has the probability of $1-\alpha$ to be the true label $Y_{test}^{*}$.
Appendix B Proofs
B.1 Proof of equation 3
@@ B.2 Proof of equation 6
Appendix C Implementation Specifics
C.1 Semantic Clustering Process
@@ Appendix E Hyperparameters
Appendix F Evaluation Metrics
@@
Appendix G Additional Experimental Analysis
@@
Appendix H Generalizability of TRON
As mentioned before, TRON applies to any generative language models that support sampling (or repeatedly responding to the same query) on various open-ended tasks. To evaluate its generalizability, we employ TRON on the (1) open-book conversational QA dataset, CoQA [Reddy et al., 2019], and the (2) closed-book reading comprehension dataset, TriviaQA [Joshi et al., 2017], utilizing LLMs, LLaMA-3-70B-Instruct [AI@Meta, 2024] and LLaMA-3.1-70B-Instruct. Note that we use LangChain with DeepInfra for LLaMA-3-70B-Instruct  [^5] https://deepinfra.com/meta-llama/Meta-Llama-3-70B-Instruct and LLaMA-3.1-70B-Instruct  [^6] https://deepinfra.com/meta-llama/Meta-Llama-3.1-70B-Instruct. We randomly select 2,000 questions from the development split of CoQA and 2,000 questions from the training split of TriviaQA, and set the split ratio between the calibration and test set to 0.5 for both datasets. Empirical results of the correctness miscoverage rate, obtained from 5 trials with randomly allocated calibration and test data at various user-specified risk levels are shown in Figure 8.
Figure 8: Results of the EER metric at various risk levels on the CoQA and TriviaQA tasks utilizing LLaMA-3-70B-Instruct and LLaMA-3.1-70B-Instruct. Note that we fix the risk level $\alpha$ to 0.1.
We also utilize the PandaGPT-13B model and GPT-4o  [^7] https://platform.openai.com/docs/models/gpt-4o on the Visual Question Answering (VQA) dataset [Antol et al., 2015] for image understanding tasks. We randomly select 1,200 VQA samples and employ 200 samples as calibration data and 1,000 samples as test data. We set $\alpha$ to 0.1 and evaluate the ERR metric at various user-specified risk levels, $\beta$, as shown in Figure 9.
Figure 9: Results of the EER metric, obtained from 5 trials with randomly allocated calibration and test data at various risk levels on the VQA dataset utilizing PandaGPT-13B and GPT-4o.
Empirical results demonstrate that TRON can also provide risk management for both MLLMs and LLMs on image understanding, conversational QA, and reading comprehension tasks. Note that while the theoretical guarantee of correctness coverage is rigorous, there can be minor fluctuations of empirical miscoverage rates that exceed the user-specified risk levels in practice due to finite-sample variability [Angelopoulos and Bates, 2021, Ye et al., 2024].
Comparisons on MCQA tasks.
We also compare TRON with the existing method for LLMs on MCQA tasks, termed Least Ambiguous set-valued Classifiers (LAC[Sadinle et al., 2019, Ye et al., 2024], which has been proven to produce prediction sets with the smallest average size [Sadinle et al., 2019]. LAC defines the nonconformal score function as one minus the softmax score corresponding to the true label. We adopt MMLU [Hendrycks et al., 2020] as the evaluation dataset utilizing LLaMA-3-8B-Instruct  [^8] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct and LLaMA-3.1-8B-Instruct  [^9] https://huggingface.co/meta-llama/Meta-Llama-3.1-8B-Instruct. To maintain consistency with previous work settings [Ye et al., 2024], we only select samples that cover at least one correct response within the candidate set when $M$ is set to 5 (i.e., $\alpha=0$). We set the splitting ratio between calibration and test set to 0.5 and $\beta$ to 0.1, and evaluate the APSS metric. Table 3 demonstrates that when sampling 10 times utilizing the LLaMA-3-8B-Instruct model, TRON’s performance can already match that of LAC, and when $M$ is 20, TRON is more predictive efficient than LAC. Additionally, when the sampling size is set to 10, TRON outperforms LAC. Our insight is that since model internal logits can induce miscalibrated biases due to arbitrarily overfit or otherwise untrustworthy issues, we analyze the output distribution of the model through sampling, which is more sound.
Model LAC TRON ($M=5$) TRON ($M=10$) TRON ($M=20$)
LLaMA-3-8B-Instruct $2.93_{(2)}$ $3.06_{(0)}$ $2.93_{(7)}$ $\mathbf{2.76_{(3)}}$
LLaMA-3.1-8B-Instruct $2.57_{(8)}$ $2.61_{(3)}$ $2.53_{(6)}$ $\mathbf{2.50_{(4)}}$
Table 3: The APSS metric ($\downarrow$) on the MMLU dataset.